Integrate the following function w.r.t x:
a) X3 + √x
b)(sinX - 2cosX)
1
Expert's answer
2020-07-27T09:34:35-0400

(a)

"\int (x^3+\sqrt{x})dx=\frac{x^4}{4}+\frac{2}{3}\sqrt{x^3}+C"

(b)

"\int (\sin{x}-2\cos{x})dx=-\cos{x}-2\sin{x}+C"